This is probability :)
Vanessa has a box of grease pencils. 10 are green, 8 are yellow, 5 are brown, and 9 are blue. If Desirre selects a grease pencil at random, which color is she twice as likely to select as brown?

Answer by rothauserc(4718)

There are 32 grease pens, the probabilities (Pr) are
:
Pr(g) = 10/32
:
Pr(y) = 8/32
:
Pr(br) = 5/32
:
Pr(bl) = 9/32
:
The Pr for selecting a green grease pen is twice that of a brown